Consider the following information: s = 0.38, dM/dY = 0.12; dY/dM = 0.35, dG = 120...

  1. Consider the following information: s = 0.38, dM/dY = 0.12; dY/dM = 0.35, dG = 120 (millions of dollars) where s = marginal propensity to save dM = change in imports, dY = change in income and dG = change in government expenditures (millions of dollars). Compute the change in income due to the government expenditure multiplier (Show the Computation and formula!)
Answer #1

Multiplier=1/(mps+mpm ) { in absence of variable tax

Mps=0.38

MPM=0.12

So ,

Multiplier=1/(0.38+0.12)=1/0.5=2

Increase in Income due to increase in government spending=120*2=240